2. Given S(x)= vand g(x) = - , find: a) (808)(*) b) What is the domain of (8°f)(x)? c) (gºf)(4)
show work please
Answer #1

Scle F(X) = 960) = -36 gof)(x) = g(60X) a) X-6 gpcx) = -3(x-6) As @ Domain of guf(x) is all those - points where denominator